(spte) 2. A Gallup Poll in June 2004 asked U.S. adults how likely they were to read Bill Clintons auto My Life. Heres how they responded Number 90 211 322 382 Response Will definitely read it Will probably read it Will probably not read it Will definitely not read it Use this information to approximate the following probabilities of a randomly selected person that participated in the poll. (a) (2 pts) What is the Pr(Will definitely read it)? (b) (2pts) What is the Pr(Will probably not read it)? (c) (2pts) What is the Pr(Will probably not read it n Will probably read it)? (d) (2pts) What is the Pr(Will probably not read it Will probably read it)?
Answer #1

total surveyed people=90+211+322+382=1005

a)Pr(will definitely read it)=90/1005=0.0896

b)P(will not probably read it )c =1-P(will not probably read it ) =1-322/1005=0.6796

c)P(will probably not read n will probably read it )=0 (As events are mutually exclusive)

d)Probability =1-P(probably not read it or probably read it)=1-(211+322)/1005=0.4697